OpenStudy (anonymous):

What is the height of a triangle with a base of 7 feet and an area of 31.5 square feet?

OpenStudy (campbell_st):

well the area formula is \[A = \frac{1}{2} \times base \times height\] so substituting you know \[31.5 = \frac{1}{2} \times 7 \times height\] solve for height

OpenStudy (anonymous):

9